OpenStudy (anonymous):

can anyone solve this find the first 5 terms of the sequence a1=500, an=(an-1)/5

OpenStudy (ash2326):

a1= 500 an= (a(n-1))/5 so a2= a(2-1)/5 a2= a1/5 a1=500 a2=500/5=100 similarly a3=a2/5=100/5=20 now the remaining two terms, you try to find out

OpenStudy (anonymous):

thank you for helping

OpenStudy (ash2326):

Glad to help :)